OpenStudy (anonymous):

Medal and fan please help :( An expression is shown below: f(x) = -16x2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

OpenStudy (anonymous):

@Hotchellerae21 could you please help me? I have confused myself so bad.

OpenStudy (anonymous):

@bohotness areyou available?

OpenStudy (anonymous):

I have been on this forever and have no idea what to do :(

OpenStudy (bohotness):

sorry i am lagg

OpenStudy (anonymous):

its okay, do you know how I could do this problem?

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20140321095831AA3ra6N its not the same one but it shows how to solve the problem

OpenStudy (anonymous):

what...

OpenStudy (lalipapa123):

can i answer for this question!!!

OpenStudy (bohotness):

go head lol

OpenStudy (anonymous):

Yes please @lalipapa123

OpenStudy (anonymous):

-16x2 + 60x + 16 = 0 16 + 60x + -16x2 = 0 16 + 60x + -16x2 = 0 4(4 + 15x + -4x2) = 0 4((4 + -1x)(1 + 4x)) = 0 4 + -1x = 0

OpenStudy (anonymous):

is what I got so far...

OpenStudy (anonymous):

@bohotness is that right for part a?

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

@iGreen could you help me?

OpenStudy (anonymous):

Its completely right?

OpenStudy (anonymous):

@Preetha would you be so kind as to help me?

OpenStudy (anonymous):

@iambatman are you available?

OpenStudy (anonymous):

Forget it. I have asked probably 10 questions in the last 2 weeks and no one can give me any explanation.